4. Sketch the domain of integration of I = order of integration. VT dy dr. Evaluate I by changing the (20p)
Answer #1

I=\int_{0}^{1}\int_{\sqrt[3]{x}}^{1}\frac{1}{y^4+1}dydx

ΟΣ

ISA VI 2

C

h 5 xso

0\leq y \leq 1

I=\int_{0}^{1}\int_{0}^{y^3}\frac{1}{y^4+1}dxdy

=\int_{0}^{1}[x]_{0}^{y^3}\frac{1}{y^4+1}dy

=\int_{0}^{1}[y^3-0]\frac{1}{y^4+1}dy

=\int_{0}^{1}\frac{y^3}{y^4+1}dy

Putting

y^4+1=t

4y^3dy=dt

y^3dy=dt/4

=\int_{1}^{2}\frac{1}{t}(dt/4)

=\frac{1}{4}\int_{1}^{2}\frac{1}{t}dt

=\frac{1}{4} [\ln t ]_{1}^{2}

=\frac{1}{4} [\ln 2-\ln 1 ]

\boldsymbol{=\frac{\ln 2}{4} }